OpenStudy (anonymous):

The x-intercept of the graph of ƒ(x)=3log(x+5)+2 is: A. 10^(2/3)-5 B. 10^(-2/3)+5 C. 10^(-2/3)-5 D. 10^(2/3)+5

OpenStudy (anonymous):

The x intercept is where y=0 0=3log(x+5)+2 -2/3=log(x+5) \[10^{\frac{-2}{3}}=x+5 \rightarrow x=10^{\frac{-2}{3}}-5\approx -4.78\]

OpenStudy (anonymous):

Choice C.

OpenStudy (anonymous):

thanks!